How much charge passes through a starting motor, if it takes 4.0 s to start a car and there is a current of 255 A during that time?
1
Expert's answer
2012-03-30T06:59:09-0400
The coulomb is defined as the charge transported by a steady current of one ampere in one second. So,

Q = I·A = 4·255& = 1020 C.